OpenStudy (anonymous):

Using the following equation, find the center and radius: x2 + 2x + y2 + 4y = 20 The center is located at (1, 2), and the radius is 25. The center is located at (−1, 2), and the radius is 25. The center is located at (−1, −2), and the radius is 5. The center is located at (1, 2), and the radius is 5.

OpenStudy (anonymous):

@emcrazy14

OpenStudy (anonymous):

@AravindG @agent0smith @B.Sanders @Bsho1997

OpenStudy (anonymous):

the center of what?

OpenStudy (anonymous):

its to find the center of a circle

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so do you know how i solve this

OpenStudy (anonymous):

x^2 + 2x + y^2 + 4y = 20 (x^2 + 2x) + (y^2 + 4y) = 20 (x^2 + 2x + 1 - 1) + (y^2 + 4y + 4 - 4) = 20 ((x^2 + 2x + 1) - 1) + ((y^2 + 4y + 4) - 4) = 20 ((x+1)^2 - 1) + ((y+2)^2 - 4) = 20 (x+1)^2 - 1 + (y+2)^2 - 4 = 20 (x+1)^2 + (y+2)^2 - 5 = 20 (x+1)^2 + (y+2)^2 = 20 + 5 (x+1)^2 + (y+2)^2 = 25 (x-(-1))^2 + (y-(-2))^2 = 25 This equation is in the form (x-h)^2 + (y-k)^2 = r^2 where (h,k) = (-1,-2) and r = 5 This is a circle with "The center is located at (-1, -2), and the radius is 5" (which is chioce c)

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

your welcome

OpenStudy (anonymous):

A circle is represented by the equation below: (x + 8)2 + (y − 3)2 = 100 Which statement is true?

OpenStudy (anonymous):

The circle is centered at (−8, 3) and has a radius of 20. The circle is centered at (8, −3) and has a diameter of 20. The circle is centered at (8, −3) and has a radius of 20. The circle is centered at (−8, 3) and has a diameter of 20.

OpenStudy (anonymous):

well using the equation of a circle which is (x−h)^2+(y−k)^2=r^2 so your center (-8,3) and your radius would be 10 the diameter for a circle is the radius times 2 so the diameter would be 20

OpenStudy (anonymous):

okay so D

OpenStudy (anonymous):

thats what i just got when I solved it

OpenStudy (anonymous):

lol yup yup